Reports 1-1 of 1 Clear search Modify search
Environmental Monitoring (Environmental Monitoring)
fiori, paoletti - 18:28 Tuesday 24 March 2020 (48796) Print this report
ENV activity of this morning

This morning we setup some monitors and injection tools which can allow some noise hunting from remote (if possible). In particular we:

  • placed two magnetic probes inside SQZ racks (detection EE room), names are: ENV_DER_MAG_1, ENV_DER_MAG_2 (I let Federico adding the pictures he took)
  • installed the small shaker (the big one resulted damaged after the last injections) clamped at the feet of DT cryotrap, it is connected to DAC channel in TCSroom (pictures attached); we tested a few low frequency injections with lines and colored noise (Figures 3, 4 and 5) which managed to exceed by x10 accelerometers arounnd the CT (Figures 3, 4, 5).
  • installed the louspeaker setup (thanks to Riccardo Romboli) in the CEB North terrace connecting it to DAC channel n.7 in DETterrace. Even in standby state the loudspekers were emitting some noisy sound probably a mains pick up due to uneven electrical ground between the DAC and the speakers power supply. Thus, we left the speakers unplugged.We tested the injection of pair of lines 181 and 181.2Hz, Figures 6, 7.
  • The ACL of tested levels are reported in the attached log.file.
Images attached to this report
Non-image files attached to this report
Comments to this report:
fiori - 16:36 Thursday 26 March 2020 (48819) Print this report

today i moved the two magnetic probes:

  • ENV_DER_MAG_1 now as shown in picture 1
  • ENV_DER_MAG_2 now as shown in picture 2
Images attached to this comment
Paolett, Fiori - 11:45 Friday 27 March 2020 (48825) Print this report

Both magnetic probes have interesting structures, varying with time (see attached, first position and second position)

However none of these is visible in Hrec, and also the coherence does not show any relevant peak.

Images attached to this comment
Search Help